OpenStudy (anonymous):

how to solve y-4/5=2 1/2

OpenStudy (anonymous):

multiply everything by 10

OpenStudy (anonymous):

how that makes no sense to me ive tried everything been stuck on this prob for forever can you please just answer it

OpenStudy (anonymous):

wat are the options

OpenStudy (anonymous):

could some one just solve this problem

OpenStudy (anonymous):

\[y-(4/5)=2+(1/2) or (y-4)/5=2+(1/2)?\]

OpenStudy (anonymous):

@shantallyd

OpenStudy (anonymous):

the first one except its 2 and 1/2 like a mixed number

OpenStudy (anonymous):

(The equation thing doesn't have mixed numbers so I'll represent 2 and 1/2 as the simple fraction 5/2 (but 2 and 1/2=2+1/2, that's what mixed numbers mean)). \[y-(4/5)=5/2\]Add 4/5 to both sides:\[y=5/2+4/5\]Bring to LCD:\[y=(5/2)*(5/5)+(4/5)*(2/2)\]\[y=(25/10)+(8/10)\]Simplify:\[y=33/10\]This is the simplest form as 33 (3*11) shares no factors with 10 (2*5), but the mixed number form is 3 + 3/10.

OpenStudy (anonymous):

3 and 3/10, if you want.

OpenStudy (anonymous):

finally thanx

OpenStudy (anonymous):

Alternatively, multiply everything by 10: \[10*(y-(4/5))=10*(5/2)\]By distributive law:\[10*y-10*(4/5)=10*(5/2)\]\[10y-40/5=50/2\]\[10y-8=25\]Add 8:\[10y=33\]Divide by 10:\[y=33/10\]It's the same both ways.